# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 2.1. Assumptions
For all $i\in\mathcal{N}$, $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ is a $\mathcal{K}_{i}$-convex function that is differentiable on an open set containing $\mathop{\bf dom}\phi_{i}$. Suppose there exists a constant $L_{g_{i}}>0$ such that the Jacobian $\mathbf{J}g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}\times n}$ satisfies
@@
Assumption 2 is standard in the analysis of first-order algorithms for distributed optimization, and has been adopted in many works (e.g., [Scaman2017, Jakovetic2018, Pu2020]). Assumption 3 has also commonly appeared in the literature related to constrained optimization problems with functional constraints [hamedani2021primal, aybat2019distributed, hamedani2021decentralized]. If the $\mathop{\bf dom}\phi$ is bounded and the Jacobian is continuous on the closure of the domain, it will be satisfied. Assumption 4 requires existence of a primal-dual solution to (1.1), which is guaranteed to hold under some mild regularity conditions, e.g., whenever (1.1) admits a Slater point. Moreover, when $g_{i}(\cdot)$ is affine, i.e., when $g_{i}(x)=A_{i}x+b_{i}$ for some $A_{i}\in\mathbb{R}^{m_{i}\times n}$ and $b_{i}\in\mathbb{R}^{m_{i}}$, the knowledge of a dual bound $B_{i}$ for $g_{i}(x)\in\mathcal{K}_{i}$ is not required. It is essential to emphasize that due to conic structure of the constraints, one can still consider a rich class of nonlinear constraints even when $g_{i}$ is affine, one can model convex quadratic inequality constraints, SOCP and SDP constraints. Furthermore, in the following remark, we discuss how the agents can compute dual bounds $B_{i}$ (as defined in Assumption 4) for $\theta_{i}^{*}$ corresponding to the constraints with nonlinear $g_{i}$ in the distributed computation setting we assume in this paper.
@@
Let $\mathcal{K}=\mathbb{R}^{N}_{+}$ and $\phi_{i}(x)=\text{1}_{X}(x)$ for $i\in\mathcal{N}$, where $X=\{x\in\mathbb{R}^{n}:\ \|x\|\leq r\}$. Consider $(P):\ \varphi^{*}=\min_{x}\{\sum_{i\in\mathcal{N}}f_{i}(x):\ g_{i}(x)\leq 0,\ i\in\mathcal{N},\ x\in X\}$. In this remark, we discuss how one can compute a Slater point for $(P)$ in a distributed manner.
@@
$({\rm Phase~I}):\quad(x^{\circ},t^{\circ})\in$ $\operatorname*{\mathbf{argmin}}_{x,t}\ \text{1}_{X^{\circ}}(x)+t$ (2.2)
@@
$\bar{g}-\underline{g}+1\geq\sum_{i\in\mathcal{N}}\theta_{i}^{\circ}\big(\bar{t}-g_{i}(\bar{x})\big)\geq\sum_{i\in\mathcal{N}}\theta_{i}^{\circ};$ (2.3)
@@ 2.2. Main Results
In this paper, we propose D-APDB, displayed in algorithm 1, for solving the constrained, composite convex consensus optimization problem in (1.1), and propose D-APDB0, displayed in algorithm 2, for solving the unconstrained [^2] Here “unconstrained” means $g_{i}(\cdot)=0$ for all $i\in\mathcal{N}$. Any node $i\in\mathcal{N}$ can still have a simple set constraint incorporated in (1.1) through choosing $\phi_{i}$ as an indicator function. version of the composite convex consensus optimization problem. Given the parameters $\alpha_{i}^{k},\beta_{i}^{k}$ and $\tilde{\alpha}_{i}^{k+1},\tilde{\beta}_{i}^{k+1},\tilde{\varsigma}_{i}^{k+1}$ as set in D-APDB, the backtracking test function $E_{i}^{k}(\cdot,\cdot)$ in Line 18 of D-APDB is defined as follows:
@@
where $\Lambda_{i}(x)\triangleq f_{i}(x)-f_{i}(x_{i}^{k})-\left\langle\nabla f_{i}(x_{i}^{k}),~x-x_{i}^{k}\right\rangle$, for all $i\in\mathcal{N}$.
@@
Suppose Assumptions 2, 3, 1 and 4 hold, and $\delta,c_{\alpha},c_{\beta},c_{\varsigma}>0$ are given such that $\delta+c<1$, where $c\triangleq c_{\alpha}+c_{\beta}+c_{\varsigma}$. Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathcal{K}^{*}$ denote an arbitrary primal-dual optimal pair satisfying Assumption 4 and $\lambda^{*}\in\mathbb{R}^{n|\mathcal{E}|}$ be an optimal dual variable corresponding to the consensus constraint $A\mathbf{x}=0$ in (2.1). For all $i\in\mathcal{N}$, it holds for all step size parameter values $\bar{\tau}_{i},\zeta_{i}>0$ that when initialized from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}=0_{m_{i}}$, the ergodic iterate sequences $\{(\bar{x}_{i}^{k},\bar{\theta}_{i}^{k})\}_{k\geq 0}$ generated by D-APDB, displayed in algorithm 1, satisfy
@@
The proof of this result given in Section 6 in three parts. The first part is on establishing the rate result which is stated in more detail with explicit $\mathcal{O}(1)$ constants and proved in corollary 6.5. In the second part, we provide an upper bound on the total number of gradient and projection evaluations required for each $i\in\mathcal{N}$ to collectively compute an $\epsilon$-optimal solution as in (6.25). Finally, in Theorem 6.7, we show that the iterate sequence converges to a primal-dual optimal solution of to (1.1).
@@
2:Inputs for each $i\in\mathcal{N}$: initial states $x_{i}^{0}\in\mathop{\bf dom}\phi_{i},\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$; step size parameters ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\tau}_{i}},\zeta_{i}>0$
@@
14:     $\tilde{\alpha}_{i}^{k+1}\leftarrow c_{\alpha}/\tilde{\tau}_{i}^{k},\quad\tilde{\beta}_{i}^{k+1}\leftarrow c_{\beta}/\tilde{\tau}_{i}^{k},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\varsigma}_{i}^{k+1}}\leftarrow c_{\varsigma}/\tilde{\tau}_{i}^{k}$
@@
18:     if $E_{i}^{k}(\tilde{x}_{i}^{k+1},\tilde{\theta}_{i}^{k+1})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{\delta}{\tilde{\tau}_{i}^{k}}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2}-\frac{\delta}{\tilde{\sigma}_{i}^{k}}\|\tilde{\theta}_{i}^{k+1}-\theta_{i}^{k}\|^{2}}$ then
@@
39:    $\alpha_{i}^{k+1}\leftarrow c_{\alpha}/\tau_{i}^{k},\quad\beta_{i}^{k+1}\leftarrow c_{\beta}/\tau_{i}^{k},\quad\varsigma_{i}^{k+1}\leftarrow c_{\varsigma}/\tau_{i}^{k}$
@@ 3. Motivation and Design of the Proposed Algorithms: D-APDB and D-APDB0
where $h:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ denotes the indicator function of $\mathcal{K}^{*}\cap\mathcal{B}$, i.e., $h(\theta)=0$ if $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$, and $\infty$ otherwise, i.e., $h(\theta)=\sum_{i\in\mathcal{N}}\text{1}_{\mathcal{K}_{i}^{*}\cap\mathcal{B}_{i}}(\theta_{i})$.
@@ 3.1. Preliminaries
We briefly present a previous work  [hamedani2021primal], where we proposed an accelerated primal-dual (APD) algorithm for solving convex-concave saddle-point (SP) problems. As it is shown in [hamedani2021primal], APD can be viewed as an extension of the primal-dual algorithm proposed in [chambolle2016ergodic] from bilinear SP problems to a more general setting with a non-bilinear coupling term. Let $\mathcal{X}\subset\mathbb{R}^{n_{x}}$ and $\mathcal{Y}\subset\mathbb{R}^{n_{y}}$ be finite-dimensional vector spaces. Here we present a slightly modified version of APD proposed in [hamedani2021decentralized] to solve the following problem:
@@
where $\|\mathbf{x}\|_{Q_{x}}=\sqrt{\mathbf{x}^{\top}Q_{x}\mathbf{x}}$ for $\mathbf{x}\in\mathcal{X}$ and $\|\mathbf{y}\|_{Q_{y}}=\sqrt{\mathbf{y}^{\top}Q_{y}\mathbf{y}}$ for $\mathbf{y}\in\mathcal{Y}$. Based on the discussion in [hamedani2021primal], if $Q_{x}$ and $Q_{y}$ are chosen such that there exist some $c\geq 1$, and $\alpha,\beta\geq 0$ satisfying
@@
– see [hamedani2021primal, Assumption 3] with $\delta=0$, then according to [hamedani2021primal, Eq. (3.3)] and [hamedani2021primal, Lemma 3.4], it holds for all $K\geq 0$ that
@@
The first goal is mainly motivated by the primal-dual dynamics for the particular saddle point formulation we focus on. Indeed, $\mathcal{H}(\mathbf{x},\mathbf{y})$ has a curvature in $\mathbf{x}$, while it is affine in $\mathbf{y}$, and that is why using momentum in $\mathbf{x}$-updates helps dampening the oscillatory behavior that naturally arises due to primal-descent-dual-ascent-type updates, e.g., see [schafer2019competitive, zheng2024dissipative] for the oscillatory behavior of gradient-descent-ascent updates; on the other hand, the use of momentum in $\mathbf{y}$-update as in (3.3) may even exacerbate the oscillations as $\nabla_{\mathbf{y}}\mathcal{H}(\mathbf{x},\mathbf{y})$ only depends on $\mathbf{x}$, i.e., if $\{\mathbf{x}^{k}\}$ oscillates, not only there is no damping mechanism for $\mathbf{x}$-updates in (3.3), but oscillations in $\{\mathbf{x}^{k}\}$ also causes $\{\mathbf{y}^{k}\}$ to oscillate even more.
@@
with $Q_{x}^{k}=Q_{x}$, $Q_{y}^{k}=Q_{y}$, and $\eta^{k}=1$ for all $k\geq 0$. A convergence rate result similar to (3.5) continue to hold for this variant with $Q_{x}$ and $Q_{y}$ chosen slightly different than (3.4) due to primal-dual switch:
$Q_{y}\;\succeq\;L_{yy}\,I_{n_{y}}\;+\;\frac{1}{\alpha}\,L_{xy}^{\top}L_{xy},\qquad Q_{x}\;-\;L_{q_{x}}\;\succeq\;c\,(\alpha+\beta)\,I_{n_{x}}\;+\;\frac{1}{\beta}\,L_{xx}^{2}\,I_{n_{x}},$ (3.7)
for some $c\geq 1$ and $\alpha,\beta\geq 0$. Our motivation for giving eq. 3.6 with iteration dependent step-size matrices $Q_{x}^{k},Q_{y}^{k}$ and momentum parameter $\eta^{k}$ is to set the grounds for a setting that will allow us to analyze a variant of APD method employing node-specific local step-size search. In the rest, we use the notation $\mathbf{x}^{k}=[x_{i}^{k}]_{i\in\mathcal{N}}$ for $k\geq 0$.
Towards the second goal stated above, i.e., to ensure that the nodes can adaptively select their primal and dual step sizes at each iteration through backtracking, we still need to slightly modify the updates given (3.6) since the $\lambda$-update in eq. 3.6 related to the dual variable vector $\lambda$ corresponding to the consensus constraint $A\mathbf{x}=0$ is not suitable for a distributed local step size search mechanism. Indeed, given some $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$ and $\gamma^{k}>0$ for $k\geq 0$, let
$Q_{x}^{k}=\mathbf{D}_{\tau^{k}},\qquad Q_{y}^{k}=\begin{pmatrix}\mathbf{D}_{\sigma^{k}}&0\\ 0&\frac{1}{\gamma^{k}}I_{n|\mathcal{E}|}\end{pmatrix}$, (3.8)
where $\mathbf{D}_{\tau^{k}}=\mathop{\bf diag}\left(\left[\frac{1}{\tau_{i}^{k}}\;I_{n}\right]_{i\in\mathcal{N}}\right)$, $\mathbf{D}_{\sigma^{k}}=\mathop{\bf diag}\left(\left[\frac{1}{\sigma_{i}^{k}}\;I_{m_{i}}\right]_{i\in\mathcal{N}}\right)$. For any iteration $k\geq 0$, after a candidate point $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$ for the next iteration is computed according to eq. 3.6 using some given step sizes $\{\tau_{i}^{k},\sigma_{i}^{k}\}_{i\in\mathcal{N}}$ and $\gamma^{k}$, one needs to check a backtracking condition to ensure that $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$ provides a sufficient decrease in a suitably defined potential function, where $\mathbf{y}^{k+1}=(\theta^{k+1},\lambda^{k+1})$ –if the condition holds, we accept $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$; otherwise, we decrease the given step sizes and compute a new candidate point. In this procedure, one needs to compute a term involving the candidate dual point $\lambda^{k+1}$ every time the condition is checked, e.g., $\|\lambda^{k+1}-\lambda^{k}\|^{2}$. However, the update rule in (3.6) implies that $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A\mathbf{x}^{k+1}$, which is a function of $\mathbf{x}^{k+1}$ for $k\geq 0$; hence, checking a test function involving $\lambda^{k+1}$ is not suitable for distributed computation. Indeed, as an example, consider the quantity $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ which can be computed as below:
@@
According to eq. 3.9, at iteration $k\geq 0$, for each agent $i\in\mathcal{N}$, checking its local backtracking condition involves computing the term $\langle x_{i}^{k+1},\sum_{j\in\mathcal{N}_{i}}(x_{i}^{k+1}-x_{j}^{k+1})\rangle$, which requires the decision vectors $x_{j}^{k+1}$ of its neighbors $j\in\mathcal{N}_{i}$ that satisfy their own backtracking conditions. Therefore, if one adopts the update scheme in (3.6), it is not trivial or practical to compute $\{x_{i}^{k+1}\}_{j\in\mathcal{N}}$ in a distributed manner that satisfy all local backtracking conditions simultaneously. To avoid this issue, we slightly change the update order as stated in eq. 3.10, i.e., at iteration $k\geq 0$, consensus related dual variable $\lambda$ is updated first using the information related to $\mathbf{x}^{k}$ and $\mathbf{x}^{k-1}$, then the $\mathbf{x}$-update is executed, and finally the dual variable $\theta_{i}$ related to the local constraint $g_{i}(x_{i})\in\mathcal{K}_{i}$ is updated for all $i\in\mathcal{N}$. Hence, for the variant of APD in eq. 3.10, the term $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ only relies on the points $\{x_{i}^{k}\}_{i\in\mathcal{N}}$ that satisfy the local backtracking conditions at iteration k1k-1. It should be emphasized that as an artifact of the proof technique we used, rather than setting $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A\mathbf{x}^{k}$, we also use momentum term in $\lambda$-update as well, i.e., $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A((1+\eta^{k})\mathbf{x}^{k}-\eta^{k}\mathbf{x}^{k-1})$ for some properly chosen $\eta^{k}\geq 1$. [^3] Later in Lemma 4.1, we establish how duality gap changes after one iteration of D-APDB. This choice of $\lambda$-update leads to a telescoping sum for the duality gap bound in (4.5) by ensuring that $t_{k+1}Q^{k+1}(\mathbf{z})-t_{k}R^{k+1}(\mathbf{z})\leq 0$ holds for all $k\geq 0$ and $\mathbf{z}\in\mathcal{X}\times\mathcal{Y}$.

### v2

@@ 2.1. Assumptions
For all $i\in\mathcal{N}$, $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ is a $\mathcal{K}_{i}$-convex function that is differentiable on an open set containing $\mathop{\bf dom}\phi_{i}$. Suppose there exists a constant $L_{g_{i}}\geq 0$ such that the Jacobian $\mathbf{J}g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}\times n}$ satisfies
@@
Assumption 2 is standard in the analysis of first-order algorithms for distributed optimization, and has been adopted in many works (e.g., [21, 41, 47]). Assumption 3 has also commonly appeared in the literature related to constrained optimization problems with functional constraints [6, 19, 20]. If the $\mathop{\bf dom}\phi$ is bounded and the Jacobian is continuous on the closure of the domain, it will be satisfied. Assumption 4 requires existence of a primal-dual solution to (1.1), which is guaranteed to hold under some mild regularity conditions, e.g., whenever (1.1) admits a Slater point. Moreover, when $g_{i}(\cdot)$ is affine, i.e., when $g_{i}(x)=A_{i}x+b_{i}$ for some $A_{i}\in\mathbb{R}^{m_{i}\times n}$ and $b_{i}\in\mathbb{R}^{m_{i}}$, the knowledge of a dual bound $B_{i}$ for $g_{i}(x)\in\mathcal{K}_{i}$ is not required. It is essential to emphasize that due to conic structure of the constraints, one can still consider a rich class of nonlinear constraints even when $g_{i}$ is affine, one can model convex quadratic inequality constraints, SOCP and SDP constraints. Furthermore, in the following remark, we discuss how the agents can compute dual bounds $B_{i}$ (as defined in Assumption 4) for $\theta_{i}^{*}$ corresponding to the constraints with nonlinear $g_{i}$ in the distributed computation setting we assume in this paper.
@@
Let $\mathcal{K}=\mathbb{R}^{N}_{+}$ and $\phi_{i}(x)=\mathds{1}_{X}(x)$ for $i\in\mathcal{N}$, where $X=\{x\in\mathbb{R}^{n}:\ \|x\|\leq r\}$. Consider $(P):\ \varphi^{*}=\min_{x}\{\sum_{i\in\mathcal{N}}f_{i}(x):\ g_{i}(x)\leq 0,\ i\in\mathcal{N},\ x\in X\}$. In this remark, we discuss how one can compute a Slater point for $(P)$ in a distributed manner.
@@
$({\rm Phase~I}):\quad(x^{\circ},t^{\circ})\in$ $\operatorname*{\mathbf{argmin}}_{x,t}\mathds{1}_{X^{\circ}}(x)+t$
@@
$\bar{g}-\underline{g}+1\geq\sum_{i\in\mathcal{N}}\theta_{i}^{\circ}\big(\bar{t}-g_{i}(\bar{x})\big)\geq\sum_{i\in\mathcal{N}}\theta_{i}^{\circ};$
@@ 2.2. Main Results
In this paper, we propose D-APDB, displayed in Algorithm 1, for solving the constrained, composite convex consensus optimization problem in (1.1), and propose D-APDB0, displayed in Algorithm 2, for solving the unconstrained22Here “unconstrained” means $g_{i}(\cdot)=0$ for all $i\in\mathcal{N}$. Any node $i\in\mathcal{N}$ can still have a simple set constraint incorporated in (1.1) through choosing $\phi_{i}$ as an indicator function. version of the composite convex consensus optimization problem. Given the parameters $\alpha_{i}^{k},\beta_{i}^{k}$ and $\tilde{\alpha}_{i}^{k+1},\tilde{\beta}_{i}^{k+1},\tilde{\varsigma}_{i}^{k+1}$ as set in D-APDB, the backtracking test function $E_{i}^{k}(\cdot,\cdot)$ in Line 18 of D-APDB is defined as follows:
@@
where $\Lambda_{i}(x)\triangleq f_{i}(x)-f_{i}(x_{i}^{k})-\left\langle\nabla f_{i}(x_{i}^{k}),~x-x_{i}^{k}\right\rangle$, for all $i\in\mathcal{N}$– in case $L_{g_{i}}=0$, setting $\beta_{i}^{k}=0$ and $\tilde{\beta}_{i}^{k+1}=0,$ we adopt $0^{2}/0=0;$ hence, we set $\left\|\left(\mathbf{J}g_{i}(x)-\mathbf{J}g_{i}(x_{i}^{k})\right)^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\theta_{i}^{k}}\right\|^{2}/\tilde{\beta}_{i}^{k+1}=0.$
@@
Suppose that Assumptions 2, 3, 1 and 4 hold, and $\delta,c_{\alpha},c_{\beta},c_{\varsigma}>0$ are given such that $\delta+c<1$, where $c\triangleq c_{\alpha}+c_{\beta}+c_{\varsigma}$. Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathcal{K}^{*}$ denote an arbitrary primal-dual optimal pair satisfying Assumption 4 and $\lambda^{*}\in\mathbb{R}^{n|\mathcal{E}|}$ be an optimal dual variable corresponding to the consensus constraint $A\mathbf{x}=0$ in (2.1). For all $i\in\mathcal{N}$, it holds for all step size parameter values $\bar{\tau}_{i},\zeta_{i}>0$ that when initialized from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}=0_{m_{i}}$, the ergodic iterate sequences $\{(\bar{x}_{i}^{k},\bar{\theta}_{i}^{k})\}_{k\geq 0}$ generated by D-APDB, displayed in Algorithm 1, satisfy
@@
The proof of this result is given in Section 6 in three parts. The first part is on establishing the rate result which is stated in more detail with explicit $\mathcal{O}(1)$ constants – see Corollary 6.7. In the second part, we provide an upper bound on the total number of gradient and projection evaluations required for each $i\in\mathcal{N}$ to collectively compute an $\epsilon$-optimal solution – see Corollary 6.8. Finally, in Theorem 6.9, we show that the iterate sequence converges to a primal-dual optimal solution of to (1.1).
@@
2:Inputs for each $i\in\mathcal{N}$: initial states $x_{i}^{0}\in\mathop{\bf dom}\phi_{i},\theta_{i}^{0}\in\mathcal{K}_{i}^{*}$; step size parameters ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\tau}_{i}},\zeta_{i}>0$; dual bound $B_{i}$ if $L_{g_{i}}>0$
@@
14:     $\tilde{\alpha}_{i}^{k+1}\leftarrow c_{\alpha}/\tilde{\tau}_{i}^{k},\quad\tilde{\beta}_{i}^{k+1}\leftarrow c_{\beta}/\tilde{\tau}_{i}^{k},\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\varsigma}_{i}^{k+1}}\leftarrow c_{\varsigma}/\tilde{\tau}_{i}^{k}$ $\triangleright$ $\tilde{\beta}_{i}^{k+1}=0$ if $L_{g_{i}}=0$
@@
18:     if $E_{i}^{k}(\tilde{x}_{i}^{k+1},\tilde{\theta}_{i}^{k+1})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{\delta}{\tilde{\tau}_{i}^{k}}\|\tilde{x}_{i}^{k+1}-x_{i}^{k}\|^{2}-\frac{\delta}{\tilde{\sigma}_{i}^{k}}\|\tilde{\theta}_{i}^{k+1}-\theta_{i}^{k}\|^{2}}$ then $\triangleright$ See (2.2) for $E_{i}^{k}(\cdot,\cdot)$
@@
39:    $\alpha_{i}^{k+1}\leftarrow c_{\alpha}/\tau_{i}^{k},\quad\beta_{i}^{k+1}\leftarrow c_{\beta}/\tau_{i}^{k},\quad\varsigma_{i}^{k+1}\leftarrow c_{\varsigma}/\tau_{i}^{k}$ $\triangleright$ $\beta_{i}^{k+1}=0$ if $L_{g_{i}}=0$
@@ 3. Motivation and Design of the Proposed Algorithms: D-APDB and D-APDB0
where $h:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ denotes the indicator function of $\mathcal{K}^{*}\cap\mathcal{B}$, i.e., $h(\theta)=0$ if $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$, and $\infty$ otherwise, i.e., $h(\theta)=\sum_{i\in\mathcal{N}}\mathds{1}_{\mathcal{K}_{i}^{*}\cap\mathcal{B}_{i}}(\theta_{i})$.
@@ 3.1. Preliminaries
We briefly present a related previous work  [20], where we proposed an accelerated primal-dual (APD) algorithm for solving convex-concave saddle-point (SP) problems. As it is shown in [20], APD can be viewed as an extension of the primal-dual algorithm proposed in [11] from bilinear SP problems to a more general setting with a non-bilinear coupling term. Let $\mathcal{X}\subset\mathbb{R}^{n_{x}}$ and $\mathcal{Y}\subset\mathbb{R}^{n_{y}}$ be finite-dimensional vector spaces. Here we present a slightly modified version of APD proposed in [19] to solve the following problem:
@@
where $\|\mathbf{x}\|_{Q_{x}}=\sqrt{\mathbf{x}^{\top}Q_{x}\mathbf{x}}$ for $\mathbf{x}\in\mathcal{X}$ and $\|\mathbf{y}\|_{Q_{y}}=\sqrt{\mathbf{y}^{\top}Q_{y}\mathbf{y}}$ for $\mathbf{y}\in\mathcal{Y}$. Based on the discussion in [20], if $Q_{x}$ and $Q_{y}$ are chosen such that there exist some $c\geq 1$, and $\alpha,\beta\geq 0$ satisfying
@@
– see [20, Assumption 3] with $\delta=0$, then according to [20, Eq. (3.3)] and [20, Lemma 3.4], it holds for all $K\geq 0$ that
@@
The first goal is mainly motivated by the primal-dual dynamics for the particular saddle point formulation we focus on. Indeed, $\mathcal{H}(\mathbf{x},\mathbf{y})$ has a curvature in $\mathbf{x}$, while it is affine in $\mathbf{y}$, and that is why using momentum in $\mathbf{x}$-updates helps dampening the oscillatory behavior that naturally arises due to primal-descent-dual-ascent-type updates, e.g., see [48, 56] for the oscillatory behavior of gradient-descent-ascent updates; on the other hand, the use of momentum in $\mathbf{y}$-update as in (3.3) may even exacerbate the oscillations as $\nabla_{\mathbf{y}}\mathcal{H}(\mathbf{x},\mathbf{y})$ only depends on $\mathbf{x}$, i.e., if $\{\mathbf{x}^{k}\}$ oscillates, not only there is no damping mechanism for $\mathbf{x}$-updates in (3.3), but oscillations in $\{\mathbf{x}^{k}\}$ also causes $\{\mathbf{y}^{k}\}$ to oscillate even more.
@@
with $Q_{x}^{k}=Q_{x}$, $Q_{y}^{k}=Q_{y}$, and $\eta^{k}=1$ for all $k\geq 0$. A convergence rate result similar to (3.5) continues to hold for this variant with $Q_{x}$ and $Q_{y}$ chosen slightly different than (3.4) due to primal-dual switch:
$Q_{y}\;\succeq\;L_{yy}\,I_{n_{y}}\;+\;\frac{1}{\alpha}\,L_{xy}^{\top}L_{xy},\qquad Q_{x}\;-\;L_{q_{x}}\;\succeq\;c\,(\alpha+\beta)\,I_{n_{x}}\;+\;\frac{1}{\beta}\,L_{xx}^{2}\,I_{n_{x}},$
for some $c\geq 1$ and $\alpha,\beta\geq 0$. Our motivation for giving (3.6) with iteration dependent step-size matrices $Q_{x}^{k},Q_{y}^{k}$ and momentum parameter $\eta^{k}$ is to set the grounds for a setting that will allow us to analyze a variant of APD method employing node-specific local step-size search. In the rest, we use the notation $\mathbf{x}^{k}=[x_{i}^{k}]_{i\in\mathcal{N}}$ for $k\geq 0$.
Towards the second goal stated above, i.e., to ensure that the nodes can adaptively select their primal and dual step sizes at each iteration through backtracking, we still need to slightly modify the updates given in (3.6) since the $\lambda$-update in (3.6), related to the dual variable vector $\lambda$ corresponding to the consensus constraint $A\mathbf{x}=0$, is not suitable for a distributed local step size search mechanism. Indeed, given some $\tau_{i}^{k},\sigma_{i}^{k}>0$ for $i\in\mathcal{N}$ and $\gamma^{k}>0$ for $k\geq 0$, let
$Q_{x}^{k}=\mathbf{D}_{\tau^{k}},\qquad Q_{y}^{k}=\begin{pmatrix}\mathbf{D}_{\sigma^{k}}&0\\ 0&\frac{1}{\gamma^{k}}I_{n|\mathcal{E}|}\end{pmatrix}$,
where $\mathbf{D}_{\tau^{k}}=\mathop{\bf diag}\left(\left[\frac{1}{\tau_{i}^{k}}\;I_{n}\right]_{i\in\mathcal{N}}\right)$, $\mathbf{D}_{\sigma^{k}}=\mathop{\bf diag}\left(\left[\frac{1}{\sigma_{i}^{k}}\;I_{m_{i}}\right]_{i\in\mathcal{N}}\right)$. For any iteration $k\geq 0$, after a candidate point $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$ for the next iteration is computed according to (3.6) using some given step sizes $\{\tau_{i}^{k},\sigma_{i}^{k}\}_{i\in\mathcal{N}}$ and $\gamma^{k}$, one needs to check a backtracking condition to ensure that $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$ provides a sufficient decrease in a suitably defined potential function, where $\mathbf{y}^{k+1}=(\theta^{k+1},\lambda^{k+1})$ –if the condition holds, we accept $(\mathbf{x}^{k+1},\mathbf{y}^{k+1})$; otherwise, we decrease the given step sizes and compute a new candidate point. In this procedure, one needs to compute a term involving the candidate dual point $\lambda^{k+1}$ every time the condition is checked, e.g., $\|\lambda^{k+1}-\lambda^{k}\|^{2}$. However, the update rule in (3.6) implies that $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A\mathbf{x}^{k+1}$, which is a function of $\mathbf{x}^{k+1}$ for $k\geq 0$; hence, checking a test function involving $\lambda^{k+1}$ is not suitable for distributed computation. Indeed, as an example, we consider the quantity $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ which can be computed as below:
@@
According to (3.7), at iteration $k\geq 0$, for each agent $i\in\mathcal{N}$, checking its local backtracking condition involves computing the term $\langle x_{i}^{k+1},\sum_{j\in\mathcal{N}_{i}}(x_{i}^{k+1}-x_{j}^{k+1})\rangle$, which requires the decision vectors $x_{j}^{k+1}$ of its neighbors $j\in\mathcal{N}_{i}$ that satisfy their own backtracking conditions. Therefore, if one adopts the update scheme in (3.6), it is not trivial or practical to compute $\{x_{i}^{k+1}\}_{j\in\mathcal{N}}$ in a distributed manner that satisfy all local backtracking conditions simultaneously. To avoid this issue, we slightly change the update order as stated in (3.8), i.e., at iteration $k\geq 0$, consensus related dual variable $\lambda$ is updated first using the information related to $\mathbf{x}^{k}$ and $\mathbf{x}^{k-1}$, then the $\mathbf{x}$-update is executed, and finally the dual variable $\theta_{i}$ related to the local constraint $g_{i}(x_{i})\in\mathcal{K}_{i}$ is updated for all $i\in\mathcal{N}$. Hence, for the variant of APD in (3.8), the term $\|\lambda^{k+1}-\lambda^{k}\|^{2}$ only relies on the points $\{x_{i}^{k}\}_{i\in\mathcal{N}}$ that satisfy the local backtracking conditions at iteration k1k-1. It should be emphasized that as an artifact of the proof technique we adopted, rather than setting $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A\mathbf{x}^{k}$, we set $\lambda^{k+1}=\lambda^{k}+\gamma^{k}A((1+\eta^{k})\mathbf{x}^{k}-\eta^{k}\mathbf{x}^{k-1})$ for some properly chosen $\eta^{k}\geq 1$, i.e., we also use momentum term in $\lambda$-update as well.33Later in Lemma 4.2, we establish how duality gap changes after one iteration of D-APDB. This choice of $\lambda$-update leads to a telescoping sum for the duality gap bound in (4.4) by ensuring that $t_{k+1}Q^{k+1}(\mathbf{z})-t_{k}R^{k+1}(\mathbf{z})\leq 0$ holds for all $k\geq 0$ and $\mathbf{z}\in\mathcal{X}\times\mathcal{Y}$.